# Magnetic disk winds in protoplanetary disks:

### v2

@@ 3.1 Global picture of disk evolution
Figure 1: Evolution of the gas surface density for the protoplanetary disk models, showing the global picture. The white contours show the extent of the centrifugal disk. We note that the time intervals are not uniform and are chosen to depict viscous spread of the disks. See Table 4 for a summary of the properties of the resulting systems.
As soon as a simulation begins, the initial prestellar cloud core starts to collapse, since it is constructed to have a supercritical mass-to-flux ratio and is gravitationally unstable. The time is measured from the onset of collapse, i.e., the start of the simulation. In the initial stages of the collapse, surface density increases obeying a nearly self-similar law (Basu & Mouschovias, 1994; Basu, 1997), while the rotation profile remains rigid within about 50 au. As the cloud continues to contract, the mass of the gas in the sink cell exceeds 0.02 ${\rm M_{\odot}}$, leading to the assumption of second collapse due to dissociation of hydrogen molecule (Larson, 1969). A protostar thus forms at the coordinate center at this time, with a point-like gravitational potential. The stellar luminosity now illuminates the disk and the gas temperature in the innermost regions exceeds 600 K due to both stellar irradiation and compressional heating. The infalling gas in the innermost regions soon starts to decelerate while approaching the star and a centrifugal disk larger than the sink cell forms at about 0.022 Myr. This initial evolution before the formation of the circumstellar disk is almost identical across all simulations.
@@
The major difference between models GV-2 and GV-4 at this global scale is the extent of the disk, especially during the Class II phase. GV-2 uses the value of $\alpha_{\rm MRI}$ as $0.01$, while that for GV-4 is $10^{-4}$. We note that $\alpha_{\rm turb}$ can achieve a much lower value in the dead zone (see Eq. 10). However, outside of the dead zone ($\gtrsim 10$ au), the disk maintains the maximum possible value of $\alpha_{\rm turb}=\alpha_{\rm MRI}$. The higher viscosity in the bulk of the disk causes GV-2 to evolve on a faster viscous timescale and show a larger viscous spread. The azimuthal asymmetries and disk substructure in the outer regions also get smoothed out with higher viscosity, producing a nearly perfect axisymmetric disk. The centrifugal radius of the disk is several hundred au, which is comparable with the gas radii of some of the largest T Tauri disks such as DM Tau or IM Lup (Dartois et al., 2003; Cleeves et al., 2016). On the other hand, the disk in GV-4 shows a lesser viscous spread. Lower viscosity in the disk is unable to efficiently get rid of angular momentum and transport the gas inwards. Thus, the gas remains in the disk for a longer time and forms a smaller ($\approx 250$ au) but relatively massive disk, which shows spiral substructure even at 0.8 Myr. The disk radii and masses are discussed quantitatively in Section 3.3, with respect to synthetic observations.
@@
The next four rows show models, WI-2, WI-4 and WI-3, and WI-3a, which include the disk winds self-consistently in the evolution equations. As specified in Table 2, the first three models have an identical prescription of magnetic disk winds, while WI-3a considers adjusted wind parameters. Consider the evolution of model WI-2; the disk shows asymmetry and spirals during the initial stage of evolution. During the Class II stage, the disk is much more compact as compared to analogous gravitoviscous model GV-2 and also smaller than the low viscosity model GV-4. In the last frame, the disk radius shrinks marginally, instead of expanding, resulting in a disk radius of $\approx 150$ au. A purely viscous disk transports the angular momentum radially outwards, redistributing it within the disk, which leads to a gradual expansion. Similarly, the gravitational torques work within the disk, but unlike viscosity, gravitational forces are inherently non-local and work over larger distances via non-axisymmetric perturbations and spiral density waves (Kratter & Lodato, 2016). With winds, the disk evolution is driven by removal of angular momentum instead of its transport leading to so called “advective disk”, where the disk size tends to contract with time (Alessi & Pudritz, 2022). In our models, the exact evolution of the disk size depends on the balance between gravitoviscous forces, which work towards expanding the disk and magnetic winds, which oppose this spread.
@@
The evolution of the low and intermediate viscosity models, WI-4 and WI-3, is peculiar; we obtain centrifugal disks that are very compact. In the case of WI-4, disk extends to a maximum size of $\approx 25$ au during Class I phase, however, it soon shrinks to a much smaller size of $\approx 6$ au, which is unusually small for a Class II disk. The disk of model WI-3 achieves a maximum size of 45 au and then gradually shrinks to about $25$ au. Although disk sizes of tens of au are well within the observational expectations for a gas disk, we will later see in Section 3.3 that both models WI-4 and WI-3 produce dust disks that are unreasonably small in comparison with observations. The trend with respect to disk size across different $\alpha_{\rm MRI}$ values confirms our hypothesis of the opposing action of disk winds and gravitoviscous expansion. With the current prescription of winds, their action is significantly stronger as compared to the viscous spread produced at low $\alpha_{\rm MRI}$ values of $10^{-4}-10^{-3}$.
@@
With WI-3a, we explore the possibility of changing the wind strength by adjusting the wind-related parameters $C_{H}$ and $C_{\beta}$. For this model, a larger value of $C_{H}$ results in a more efficient mass loss. The larger value of $C_{\beta}$ used implies weaker winds, as it results in a decrease in both the wind mass loss and the surface stress (see Table 2 and Eqs. (25), (26)). As seen in the last row of Fig. 1, the disk in WI-3a achieves a larger size as compared to the other two wind models at low viscosity, while also reaching a maximum size that is smaller than the GV models. This again indicates that the effect of reducing the magnetic wind efficiency via the $C_{\beta}$ parameter is as expected; the balance shifts in favor of viscous expansion and a larger disk size is obtained. We note that the size estimates discussed here come from the theoretical centrifugal disk and thus is not directly observable. We quantitatively discuss the disk sizes and compare synthetic observations with demographics from an ALMA survey in Section 3.3. Additionally, we note that the gas surface density depicted in Fig. 1 gives us only a limited understanding of the total disk masses. For example, the disk sizes of models GV-4, WI-2 and WI-3a are comparable at any given time. However, their masses diverge significantly after the Class 0 stage. At 1 Myr, GV-4 produces a massive disk of $0.2M_{\odot}$, which is nearly twice as massive as the disk in WI-2, and over 7 times more massive than that in WI-3a. The evolution of disk masses is discussed in detail in Section 3.3 (see Fig. 5 and Table 4).
@@ 3.2 Vertical and thermo-chemical structure of the disk
In this section, we discuss the results for the physical and thermo-chemical structure of a representative disk model obtained with the FEOSAD-ProDiMo interface. Figures 2 and 3 show the model WI-3a at 0.2 and 1.0 Myr, probing the disk as it progresses through the Class II stage. The first row of Fig. 2 shows the evolving surface density profile $\Sigma_{\rm g}(r)$. The early phases are featured by a relatively flat surface density between about 1 au and 10 au, with a number of transient time-dependent pressure bumps, which are much more pronounced in dust than in gas. At later stages, only one pressure bump remains, whereas the disk outside of about 3 au follows a structure as expected from viscous spreading (see, e.g. McCaughrean & O’Dell, 1996; Hartmann et al., 1998), with an approximate radial power-law surface density profile and an exponential decay beyond a critical radius of about 50 au. Inside of about 1 au, the surface density profile increases with the radius in both gas and dust at all evolutionary phases. The overall radial structure is similar to previous investigations with FEOSAD, where concentric gas and dust rings form in the innermost $<5\,$au region of the dead zone due to viscous torques (Kadam et al., 2019, 2022). These regions are expected to host streaming instabilities and produce planetesimals (Yang et al., 2017).
A salient difference between an early (0.2 Myr) and an evolved Class II disk (1.0 Myr), according to this model, is the distribution of the dust. At 0.2 Myr, the global dust-to-gas mass ratio only marginally deviates from the initial value of 1%, with the total dust and gas masses of $9\times 10^{-4}\,M_{\odot}$ and $0.16\,M_{\odot}$, respectively. In the outer disk regions, the local dust-to-gas ratio is also maintained and is only marginally less than the initial value. However, at 1 Myr, most of the dust is found in the vicinity of the pressure maximum near 1 au. Both, the total dust and gas masses, $8.6\times 10^{-5}\,M_{\odot}$ and $0.029\,M_{\odot}$, respectively, have diminished during disk evolution approximately by a factor of 10. The outer disk regions are severely depleted of dust; the local gas-to-dust ratio beyond 3 au is $\lesssim$ 1:10000. This is a consequence of the dust dynamics via its growth in the low viscosity environment and subsequent inward radial drift.
@@
Figure 2: The disk structure of the WI-3a model after 0.2 Myr and 1 Myr. The upper plots show the gas and dust column densities, where the green dot marks the popular MMSN-value of $\rm 1700\,g/cm^{2}$ at 1 au. The second row of plots shows the scale heights $H_{g}(r)$ as passed from FEOSAD, and the values calculated from the gas temperatures calculated by ProDiMo, using either the midplane temperature only (red dashed) or the $T_{\rm gas}(z)$ structure up to a height where the radial visual extinction is $A_{\rm V,rad}\!=\!0.001$. The lower two rows show the hydrogen nuclei density $n_{\langle H\rangle}$ and the gas-to-dust mass ratio. Additional contour lines show the radial and vertical optical extinctions, $A_{\rm V,rad}$ and $A_{\rm V,ver}$, respectively.
@@
Figure 3: Continuation of Fig. 2, showing the dust temperature structure (top) and the CO-concentration (second row) calculated by ProDiMo based on the WI-3a model. Additional contour times show the ionization parameter $\chi/n$ (UV field strength divided by total particle density) to indicate where about the CO molecules photo-dissociate, and a dust temperature of 25 K to indicate where CO freezes out. The lower diagrams show the simulated 1.3 continuum and CO 2-1 line maps, from which we derive a continuum radius of 89 au and 109 au at 1.3 mm after 0.2 Myr and 1 Myr, respectively, and a disk radius in the CO 2-1 line of 198 au and 375 au, respectively. Each radius encircles 90% of the respective continuum or line flux.
@@
In the second row of Fig. 2, we compare the scale heights calculated from FEOSAD’s midplane temperature and taking disk self-gravity into account (blue), to that calculated from ProDiMo’s midplane temperature (red dashed), and a more sophisticated variant thereof (black dotted). In case of the latter, the equation of hydrostatic equilibrium is integrated upwards into the optically thin region using ProDiMo’s $T_{\rm gas}(r,z)$ and then fitted by a Gaussian of single temperature. In the early phase, the disk is taller in the vicinity of inner rings, because of the combined effect of viscous heating and large vertical dust optical depths, which hinder the viscous heat to escape and create a temperature inversion (Fig. 3). As the disk evolves, the viscous heating diminishes and the scale heights in the inner regions decrease. Hence, the disk becomes increasingly more extended and flared due to the impact of direct irradiation from the star. Since FEOSAD and ProDiMo use different opacities (see Sects. 2.1 and 2.3), it is no surprise that the independent computation of the scale heights by ProDiMo is not identical with the values passed from FEOSAD. However, the overall magnitude and shape of the calculated $H_{\!g}(r)$ matches reasonably well, especially in the outer disk beyond 10 au. Since the inner rim near the sink cell is directly illuminated by the star in ProDiMo, it is much warmer and the scale height is much larger as compared to that in FEOSAD, where this radiative effect is not included.
@@
The gas density structure shown in the third row of Fig. 2 directly follows from $\Sigma_{\rm g}(r)$ and $H_{\!g}(r)$. We obtain maximum gas particle densities of about $10^{15}\rm\,cm^{-3}$ in the pressure bumps around 1 au. The last row of Fig. 2 shows the gas-to-dust mass ratio after settling. We obtain local dust-to-gas ratios $\ga\!1$ after dust settling in the pressure bumps, which is consistent with the midplane values in FEOSAD. However, the dust-to-gas ratio is $<$ 10-4 at the disk surface where the optical radial extinction $A_{\rm V,rad}\!=\!1$, which is an important result concerning the formation of mid-IR molecular emission lines observable with Spitzer and the James Webb Space Telescope (JWST) (Woitke et al., 2024).
@@
In Fig. 3, the first row shows the dust temperature structure as calculated by ProDiMo. Due to viscous heating and close-to-diffusive radiative transfer, the midplane temperature features several smaller variations between 1 au and 5 au at an age of 0.2 Myr. Consequently, as seen in the isothermal contours, there are multiple icelines (or snow-surfaces in 2D) around 150 K behind the pressure bumps, beyond 5 au. At the later stage, the icelines sit right behind the first and only pressure bump at about 1.5 au. Considering the second row of Fig. 3, the CO molecules populate a sandwich layer between disk midplane and its surface in the outer disk beyond the pressure bumps. This narrow layer is (i) warm enough to prevent CO freeze out ($T_{\rm d}\!\ga\!25\,$K), and (ii) protected enough from UV radiation to prevent CO photo-dissociation. The latter indicated by $\chi/n\!\la\!10^{-4}$ line, where $\chi$ is the UV field strength normalize to standard solar neighborhood, and $n$ is the total hydrogen nuclei density.
@@
Table 3 shows how the derived continuum and CO line radii depend on the applied cut-off flux level, and on the choice of the line and ALMA band. The results are found to be rather independent of the choice of the ALMA band. However, the choice of flux threshold level is important, in particular for the continuum, which implies a considerable measurement uncertainty for the continuum radii. According to our model WI-3a, different threshold flux levels result in rather similar dust radii 50-70 au after 0.2 Myr, but in a large variety of values between 20 and 180 au after 1 Myr. With time, the dust grows and drifts, moving the carriers of the millimeter opacity inwards. Despite of the action of the winds, the outer disk still experiences gravitoviscous spread, which contains trace abundances of larger grains. This explains the spread in radii with threshold flux at later times and implies that observations with high signal-to-noise ratio may result in much larger continuum radii. In summary, from our reference wind model WI-3a, we conclude that the disks tend to become fainter in the continuum with age, whereas they become brighter and larger in the CO lines due to gravitoviscous spreading.

### v3

@@ 3.1 Global picture of disk evolution
Figure 1: Evolution of the gas surface density for the protoplanetary disk models showing the global picture. The white contours show the extent of the centrifugal disk. We note that the time intervals are not uniform and are chosen to depict viscous spread of the disks. (See Table 4 for a summary of the properties of the resulting systems.)
As soon as a simulation begins, the initial prestellar cloud core starts to collapse, since it is constructed to have a supercritical mass-to-flux ratio and is gravitationally unstable. The time is measured from the onset of collapse, that is, the start of the simulation. In the initial stages of the collapse, surface density increases obeying a nearly self-similar law (Basu & Mouschovias, 1994; Basu, 1997), while the rotation profile remains rigid within about 50 au. As the cloud continues to contract, the mass of the gas in the sink cell exceeds 0.02 ${\rm M_{\odot}}$, leading to the assumption of second collapse due to dissociation of hydrogen molecule (Larson, 1969). A protostar thus forms at the coordinate center at this time, with a point-like gravitational potential. The stellar luminosity now illuminates the disk and the gas temperature in the innermost regions exceeds 600 K due to both stellar irradiation and compressional heating. The infalling gas in the innermost regions soon starts to decelerate while approaching the star and a centrifugal disk larger than the sink cell forms at about 0.022 Myr. This initial evolution before the formation of the circumstellar disk is almost identical across all simulations.
@@
The major difference between models GV-2 and GV-4 at this global scale is the extent of the disk, especially during the Class II phase. GV-2 uses the value of $\alpha_{\rm MRI}$ as $0.01$, while that for GV-4 is $10^{-4}$. We note that $\alpha_{\rm turb}$ can achieve a much lower value in the dead zone (see Eq. 10). However, outside of the dead zone ($\gtrsim 10$ au), the disk maintains the maximum possible value of $\alpha_{\rm turb}=\alpha_{\rm MRI}$. The higher viscosity in the bulk of the disk causes GV-2 to evolve on a faster viscous timescale and show a larger viscous spread. The azimuthal asymmetries and disk substructure in the outer regions also get smoothed out with higher viscosity, producing a nearly perfect axisymmetric disk. The centrifugal radius of the disk is several hundred au, which is comparable with the gas radii of some of the largest T Tauri disks such as DM Tau or IM Lup (Dartois et al., 2003; Cleeves et al., 2016). On the other hand, the disk in GV-4 shows a lesser viscous spread. Lower viscosity in the disk is unable to efficiently get rid of angular momentum and transport the gas inward. Thus, the gas remains in the disk for a longer time and forms a smaller ($\approx 250$ au) but relatively massive disk, which shows spiral substructure even at 0.8 Myr. The disk radii and masses are discussed quantitatively in Sect. 3.3, with respect to synthetic observations.
@@
The next four rows show models, WI-2, WI-4 and WI-3, and WI-3a, which include the disk winds self-consistently in the evolution equations. As specified in Table 2, the first three models have an identical prescription of magnetic disk winds, while WI-3a considers adjusted wind parameters. Consider the evolution of model WI-2; the disk shows asymmetry and spirals during the initial stage of evolution. During the Class II stage, the disk is much more compact as compared to analogous gravitoviscous model GV-2 and also smaller than the low viscosity model GV-4. In the last frame, the disk radius shrinks marginally, instead of expanding, resulting in a disk radius of $\approx 150$ au. A purely viscous disk transports the angular momentum radially outward, redistributing it within the disk, which leads to a gradual expansion. Similarly, the gravitational torques work within the disk, but unlike viscosity, gravitational forces are inherently non-local and work over larger distances via non-axisymmetric perturbations and spiral density waves (Kratter & Lodato, 2016). With winds, the disk evolution is driven by removal of angular momentum instead of its transport leading to so called “advective disk”, where the disk size tends to contract with time (Alessi & Pudritz, 2022). In our models, the exact evolution of the disk size depends on the balance between gravitoviscous forces, which work toward expanding the disk and magnetic winds, which oppose this spread.
@@
The evolution of the low and intermediate viscosity models, WI-4 and WI-3, is peculiar; we obtain centrifugal disks that are very compact. In the case of WI-4, the disk extends to a maximum size of $\approx 25$ au during the Class I phase. However, it soon shrinks to a much smaller size of $\approx 6$ au, which is unusually small for a Class II disk. The disk of model WI-3 achieves a maximum size of 45 au and then gradually shrinks to about $25$ au. Although disk sizes of tens of au are well within the observational expectations for a gas disk, we later show in Sect. 3.3 that both models WI-4 and WI-3 produce dust disks that are unreasonably small in comparison with observations. The trend with respect to disk size across different $\alpha_{\rm MRI}$ values confirms our hypothesis of the opposing action of disk winds and gravitoviscous expansion. With the current prescription of winds, their action is significantly stronger as compared to the viscous spread produced at low $\alpha_{\rm MRI}$ values of $10^{-4}-10^{-3}$.
@@
With WI-3a, we explore the possibility of changing the wind strength by adjusting the wind-related parameters $C_{H}$ and $C_{\beta}$. For this model, a larger value of $C_{H}$ results in a more efficient mass loss. The larger value of $C_{\beta}$ used implies weaker winds, as it results in a decrease in both the wind mass loss and the surface stress (see Table 2 and Eqs. (25), (26)). As seen in the last row of Fig. 1, the disk in WI-3a achieves a larger size as compared to the other two wind models at low viscosity, while also reaching a maximum size that is smaller than the GV models. This again indicates that the effect of reducing the magnetic wind efficiency via the $C_{\beta}$ parameter is as expected; the balance shifts in favor of viscous expansion and a larger disk size is obtained. We note that the size estimates discussed here come from the theoretical centrifugal disk and thus are not directly observable. We quantitatively discuss the disk sizes and compare synthetic observations with demographics from an ALMA survey in Sect. 3.3. Additionally, we note that the gas surface density depicted in Fig. 1 gives us only a limited understanding of the total disk masses. For example, the disk sizes of models GV-4, WI-2 and WI-3a are comparable at any given time. However, their masses diverge significantly after the Class 0 stage. At one Myr, GV-4 produces a massive disk of $0.2M_{\odot}$, which is nearly twice as massive as the disk in WI-2, and over 7 times more massive than that in WI-3a. The evolution of disk masses is discussed in detail in Sect. 3.3 (see Fig. 5 and Table 4).
@@ 3.2 Vertical and thermo-chemical structure of the disk
In this section we discuss the results for the physical and thermo-chemical structure of a representative disk model obtained with the FEOSAD-ProDiMo interface. Figures 2 and 3 show the model WI-3a at 0.2 and one Myr, probing the disk as it progresses through the Class II stage. The first row of Fig. 2 shows the evolving surface density profile $\Sigma_{\rm g}(r)$. The early phases are featured by a relatively flat surface density between about one au and ten au, with a number of transient time-dependent pressure bumps, which are much more pronounced in dust than in gas. At later stages, only one pressure bump remains, whereas the disk outside of about three au follows a structure as expected from viscous spreading (see, e.g., McCaughrean & O’Dell, 1996; Hartmann et al., 1998), with an approximate radial power-law surface density profile and an exponential decay beyond a critical radius of about 50 au. Inside of about one au, the surface density profile increases with the radius in both gas and dust at all evolutionary phases. The overall radial structure is similar to previous investigations with FEOSAD, where concentric gas and dust rings form inside of the innermost five au region of the dead zone due to viscous torques (Kadam et al., 2019, 2022). These regions are expected to host streaming instabilities and produce planetesimals (Yang et al., 2017).
A salient difference between an early (0.2 Myr) and an evolved Class II disk (one Myr), according to this model, is the distribution of the dust. At 0.2 Myr, the global dust-to-gas mass ratio only marginally deviates from the initial value of 1%, with the total dust and gas masses of $9\times 10^{-4}\,M_{\odot}$ and $0.16\,M_{\odot}$, respectively. In the outer disk regions, the local dust-to-gas ratio is also maintained and is only marginally less than the initial value. However, at one Myr, most of the dust is found in the vicinity of the pressure maximum near 1 au. Both, the total dust and gas masses, $8.6\times 10^{-5}\,M_{\odot}$ and $0.029\,M_{\odot}$, respectively, have diminished during disk evolution approximately by a factor of 10. The outer disk regions are severely depleted of dust; the local gas-to-dust ratio beyond three au is $\lesssim$ 1:10000. This is a consequence of the dust dynamics via its growth in the low viscosity environment and subsequent inward radial drift.
@@
Figure 2: Disk structure of the WI-3a model after 0.2 Myr and one Myr. The upper plots show the gas and dust column densities, where the green dot marks the popular MMSN-value of $\rm 1700\,g/cm^{2}$ at one au. The second row of plots shows the scale heights $H_{g}(r)$ as passed from FEOSAD, and the values calculated from the gas temperatures calculated by ProDiMo, using either the midplane temperature only (red dashed) or the $T_{\rm gas}(z)$ structure up to a height where the radial visual extinction is $A_{\rm V,rad}\!=\!0.001$. The lower two rows show the hydrogen nuclei density $n_{\langle H\rangle}$ and the gas-to-dust mass ratio. Additional contour lines show the radial and vertical optical extinctions, $A_{\rm V,rad}$ and $A_{\rm V,ver}$, respectively.
@@
Figure 3: Continuation of Fig. 2 showing the dust temperature structure (top) and the CO-concentration (second row) calculated by ProDiMo based on the WI-3a model. Additional contour times show the ionization parameter $\chi/n$ (UV field strength divided by total particle density) to indicate where about the CO molecules photo-dissociate, and a dust temperature of 25 K to indicate where CO freezes out. The lower diagrams show the simulated 1.3 continuum and CO 2-1 line maps, from which we derive a continuum radius of 89 au and 109 au at 1.3 mm after 0.2 Myr and one Myr, respectively, and a disk radius in the CO 2-1 line of 198 au and 375 au, respectively. Each radius encircles 90% of the respective continuum or line flux.
@@
In the second row of Fig. 2, we compare the scale heights calculated from FEOSAD’s midplane temperature and taking disk self-gravity into account (blue), to that calculated from ProDiMo’s midplane temperature (red dashed), and a more sophisticated variant thereof (black dotted). In case of the latter, the equation of hydrostatic equilibrium is integrated upward into the optically thin region using ProDiMo’s $T_{\rm gas}(r,z)$ and then fitted by a Gaussian of single temperature. In the early phase, the disk is taller in the vicinity of inner rings, because of the combined effect of viscous heating and large vertical dust optical depths, which hinder the viscous heat to escape and create a temperature inversion (Fig. 3). As the disk evolves, the viscous heating diminishes and the scale heights in the inner regions decrease. Hence, the disk becomes increasingly more extended and flared due to the impact of direct irradiation from the star. Since FEOSAD and ProDiMo use different opacities (see Sects. 2.1 and 2.3), it is no surprise that the independent computation of the scale heights by ProDiMo is not identical with the values passed from FEOSAD. However, the overall magnitude and shape of the calculated $H_{\!g}(r)$ matches reasonably well, especially in the outer disk beyond 10 au. Since the inner rim near the sink cell is directly illuminated by the star in ProDiMo, it is much warmer and the scale height is much larger as compared to that in FEOSAD, where this radiative effect is not included.
@@
The gas density structure shown in the third row of Fig. 2 directly follows from $\Sigma_{\rm g}(r)$ and $H_{\!g}(r)$. We obtain maximum gas particle densities of about $10^{15}\rm\,cm^{-3}$ in the pressure bumps around one au. The last row of Fig. 2 shows the gas-to-dust mass ratio after settling. We obtain local dust-to-gas ratios $\ga\!1$ after dust settling in the pressure bumps, which is consistent with the midplane values in FEOSAD. However, the dust-to-gas ratio is $<$ 10-4 at the disk surface where the optical radial extinction $A_{\rm V,rad}\!=\!1$, which is an important result concerning the formation of mid-IR molecular emission lines observable with Spitzer and the James Webb Space Telescope (JWST) (Woitke et al., 2024).
@@
In Fig. 3, the first row shows the dust temperature structure as calculated by ProDiMo. Due to viscous heating and close-to-diffusive radiative transfer, the midplane temperature features several smaller variations between one and five au at an age of 0.2 Myr. Consequently, as seen in the isothermal contours, there are multiple icelines (or snow-surfaces in 2D) around 150 K behind the pressure bumps, beyond five au. At the later stage, the icelines sit right behind the first and only pressure bump at about 1.5 au. Considering the second row of Fig. 3, the CO molecules populate a sandwich layer between disk midplane and its surface in the outer disk beyond the pressure bumps. This narrow layer is (i) warm enough to prevent CO freeze-out ($T_{\rm d}\!\ga\!25\,$K), and (ii) protected enough from UV radiation to prevent CO photo-dissociation. The latter indicated by $\chi/n\!\la\!10^{-4}$ line, where $\chi$ is the UV field strength normalize to standard solar neighborhood, and $n$ is the total hydrogen nuclei density.
@@
Table 3 shows how the derived continuum and CO line radii depend on the applied cut-off flux level, and on the choice of the line and ALMA band. The results are found to be rather independent of the choice of the ALMA band. However, the choice of flux threshold level is important, in particular for the continuum, which implies a considerable measurement uncertainty for the continuum radii. According to our model WI-3a, different threshold flux levels result in rather similar dust radii 50-70 au after 0.2 Myr, but in a large variety of values between 20 and 180 au after one Myr. With time, the dust grows and drifts, moving the carriers of the millimeter opacity inward. Despite the action of the winds, the outer disk still experiences gravitoviscous spread, which contains trace abundances of larger grains. This explains the spread in radii with threshold flux at later times and implies that observations with high signal-to-noise ratio may result in much larger continuum radii. In summary, from our reference wind model WI-3a, we conclude that the disks tend to become fainter in the continuum with age, whereas they become brighter and larger in the CO lines due to gravitoviscous spreading.